You have been asked to prepare the monthly cash budget for June and July for the Merchandise and Mercantile Company. The compCr May 31 4,950 April 30 51,000 24.990 2,040 May 31 140,000 20,400 68,600 Information From Accounting Records and Planning DoRequired: 1. Calculate the cost per unit of merchandise inventory. Cost per unit 2. Prepare a schedule showing the quantity o4. Calculate the percentages of sales the company expects to collect in the month of the sale and in the two months following7. Calculate the balance in the cash account on June 30, based on the transactions projected to occur in June. Use the collec
Answer #1

Solutions:

1. June Sales = 206500

Selling price = 2065000/35 = 5900 units

Cost of Goods Sold (June) = 23600

Cost per unit = 23600/5900 = $4

2.  

May June July
Sales 4000 5900 3730
Ending Inventory 2950 1865 2365
Beginning Inventory 2000 2950 1865
Purchases 4950 4815 4230
Cost $4.00 $4.00 $4.00

Sales = (Sales Revenue / Selling Price per unit)

Ending Inventory for the current month = Sales of the following month / 2

Begining Inventory = Ending Inventory of Previous Month

Purchases = Sales +Ending Inventory - Opening Inventory

3.

May June July
Purchases $19,800.00 $19,260.00 $16,920.00

4.

% Collected In April 49.00%
% Collected in May 40.00%
% Collected in Jun 7.00%
% Uncollectible 4.00%
Total 100.00%

April = 24990/51000

May = 20400/51000

June = 3570/51000

Uncollectible = 2040/51000